# Don’t Start from Scratch: Behavioral Refinement via Interpolant-based Policy Diffusion

### v2

@@ Abstract
Imitation learning empowers artificial agents to mimic behavior by learning from demonstrations. Recently, diffusion models, which have the ability to model high-dimensional and multimodal distributions, have shown impressive performance on imitation learning tasks. These models learn to shape a policy by diffusing actions (or states) from standard Gaussian noise. However, the target policy to be learned is often significantly different from Gaussian and this mismatch can result in poor performance when using a small number of diffusion steps (to improve inference speed) and under limited data. The key idea in this work is that initiating from a more informative source than Gaussian enables diffusion methods to mitigate the above limitations. We contribute both theoretical results, a new method, and empirical findings that show the benefits of using an informative source policy. Our method, which we call BRIDGeR, leverages the stochastic interpolants framework to bridge arbitrary policies, thus enabling a flexible approach towards imitation learning. It generalizes prior work in that standard Gaussians can still be applied, but other source policies can be used if available. In experiments on challenging simulation benchmarks and on real robots, BRIDGeR outperforms state-of-the-art diffusion policies. We provide further analysis on design considerations when applying BRIDGeR.
@@ I Introduction
  • A practical method that enables source distributions to be used in diffusion methods for imitation learning, allowing for better trade-offs between inference speed and performance;
@@ II-B Diffusion-based Policy Learning
One limitation the original DDPM is slow sampling; the number of steps $K$ required is typically large (e.g., hundreds to thousands). To improve sample efficiency, diffusion-based policy methods [7, 37] employ Denoising Diffusion Implicit Models (DDIM) [9], which apply a deterministic non-Markovian process to trade-off between sample quality and inference speed.
@@ IV Method: BRIDGeR for Imitation Learning
BRIDGeR is based on Stochastic Interpolants [2], a recently-proposed framework to bridge densities in finite time. At a high-level, a stochastic interpolant is a time-dependent random variable that “interpolates” between two arbitrary densities. As a concrete example, consider two densities $p_{0}$ and $p_{1}$. A simple linear stochastic interpolant is
@@ IV-A Stochastic Interpolants for Imitation Learning
     $a_{t_{k+1}}=a_{t_{(}k)}+b_{F}(t,a_{t_{k}},x)\delta t+\sqrt{2\epsilon(t_{k})}z$
@@ V-A Domains
Door $k=0$ $0.13\pm 0.04$ $\mathbf{0.16\pm 0.06}$ $00.00\pm 0$ 0.00 $0.13\pm 0.04$ $\mathbf{0.16\pm 0.06}$ $00.00\pm 0$ 0.00
@@ VII Conclusions and Future Work
Limitations and Future Work. Here, we have shown that leveraging prior knowledge (in the form of source policies) improves learned diffusion policy performance. BRIDGeR opens up avenues to explore different forms of prior knowledge, e.g., policies designed for other tasks (transfer learning) and those constructed using foundation models. Specific to BRIDGeR, we explored a salient but limited set of design considerations; future work can examine more elaborate interpolant functions, along with more in-depth analysis of noise schedules and diffusion coefficients. Finally, we plan to extend BRIDGeR incorporate considerations such as safety constraints.
@@ Appendix F Additional Results and Plots
Door $k=0$ $0.13\pm 0.04$ $\mathbf{0.16\pm 0.06}$ $00.00\pm 0$ 0.00 $0.13\pm 0.04$ $\mathbf{0.16\pm 0.06}$ $00.00\pm 0$ 0.00
@@
Door $k=0$ $0.12\pm 0.08$ $\mathbf{0.16\pm 0.06}$ $00.00\pm 0$ 0.00 $0.12\pm 0.08$ $\mathbf{0.16\pm 0.06}$ $00.00\pm 0$ 0.00
@@
Door $k=0$ $0.08\pm 0.12$ $\mathbf{0.16\pm 0.06}$ $00.00\pm 0$ 0.00 $0.08\pm 0.12$ $\mathbf{0.16\pm 0.06}$ $00.00\pm 0$ 0.00
@@
Seen Categories $k=0$ $\mathbf{0.86\pm 0.09}$ $1.11\pm 0.06$ $1.34\pm 0.08$ $\mathbf{0.66\pm 0.09}$ $1.21\pm 0.06$ $1.34\pm 0.08$ $1.45\pm 0.08$
@@
Door $k=0$ $0.13\pm 0.04$ $\mathbf{0.16\pm 0.06}$ $00.00\pm 0$ 0.00 $0.13\pm 0.04$ $\mathbf{0.16\pm 0.06}$ $00.00\pm 0$ 0.00
@@
Door $k=0$ $0.12\pm 0.08$ $\mathbf{0.16\pm 0.06}$ $00.00\pm 0$ 0.00 $0.12\pm 0.08$ $\mathbf{0.16\pm 0.06}$ $00.00\pm 0$ 0.00
@@
Door $k=0$ $0.08\pm 0.12$ $\mathbf{0.16\pm 0.06}$ $00.00\pm 0$ 0.00 $0.08\pm 0.12$ $\mathbf{0.16\pm 0.06}$ $00.00\pm 0$ 0.00
@@
Seen Categories $k=0$ $\mathbf{0.86\pm 0.09}$ $1.11\pm 0.06$ $1.34\pm 0.08$ $\mathbf{0.66\pm 0.09}$ $1.21\pm 0.06$ $1.34\pm 0.08$

### v4

@@ Abstract
Imitation learning empowers artificial agents to mimic behavior by learning from demonstrations. Recently, diffusion models, which have the ability to model high-dimensional and multimodal distributions, have shown impressive performance on imitation learning tasks. These models learn to shape a policy by diffusing actions (or states) from standard Gaussian noise. However, the target policy to learn is often significantly different from Gaussian and this mismatch can result in poor performance when using a small number of diffusion steps (to improve inference speed) and under limited data. The key idea in this work is that initiating from a more informative source than Gaussian enables diffusion methods to mitigate the above limitations. We contribute both theoretical results, a new method, and empirical findings that show the benefits of using an informative source policy. Our method, which we call BRIDGeR, leverages the stochastic interpolants framework to bridge arbitrary policies, thus enabling a flexible approach towards imitation learning. It generalizes prior work in that standard Gaussians can still be applied, but other source policies can be used if available. In experiments on challenging simulation benchmarks and on real robots, BRIDGeR outperforms state-of-the-art diffusion policies. We provide further analysis on design considerations when applying BRIDGeR. Code for BRIDGeR is available at https://github.com/clear-nus/bridger.
@@ I Introduction
  • A practical method that enables source policies to be used for diffusion-based imitation learning, which leads to better trade-offs between inference speed and performance;
@@ II-B Diffusion-based Policy Learning
One limitation of the original DDPM is slow sampling; the number of steps $K$ required is typically large (e.g., hundreds to thousands). To improve sample efficiency, diffusion-based policy methods [7, 37] employ Denoising Diffusion Implicit Models (DDIM) [9], which apply a deterministic non-Markovian process to trade-off between sample quality and inference speed.
@@ IV Method: BRIDGeR for Imitation Learning
BRIDGeR is based on Stochastic Interpolants [2], a recently-proposed framework to bridge densities in finite time. At a high-level, a stochastic interpolant is a continuous-time stochastic process $\{y_{t}\}_{t}$ that “interpolates” between two arbitrary densities. As a concrete example, consider two densities $p_{0}$ and $p_{1}$. A simple linear stochastic interpolant is
@@ IV-A Stochastic Interpolants for Imitation Learning
     $a_{t_{k+1}}=a_{t_{k}}+b_{F}(t,a_{t_{k}},x)\delta t+\sqrt{2\epsilon(t_{k})}z$
@@ V-A Domains
Door $k=0$ $0.21\pm 0.04$ $\mathbf{0.22\pm 0.06}$ $00.00\pm 0$ 0.00 $0.21\pm 0.04$ $\mathbf{0.22\pm 0.06}$ $00.00\pm 0$ 0.00
@@ VII Conclusions and Future Work
Limitations and Future Work. Here, we have shown that leveraging prior knowledge (in the form of source policies) improves learned diffusion policy performance. BRIDGeR opens up avenues to explore different forms of prior knowledge, e.g., policies designed for other tasks (transfer learning) and those constructed using foundation models. Here, we explored a salient but limited set of design considerations; future work can examine more elaborate interpolant functions, along with more in-depth analysis of noise schedules and diffusion coefficients. Finally, we plan to extend BRIDGeR incorporate considerations such as safety and user preferences.
Acknowledgements
This research is supported by the National Research Foundation, Singapore under its Medium Sized Center for Advanced Robotics Technology Innovation.
@@ Appendix F Additional Results and Plots
Door $k=0$ $0.21\pm 0.04$ $\mathbf{0.22\pm 0.06}$ $00.00\pm 0$ 0.00 $0.21\pm 0.04$ $\mathbf{0.22\pm 0.06}$ $00.00\pm 0$ 0.00
@@
Door $k=0$ $0.20\pm 0.08$ $\mathbf{0.22\pm 0.06}$ $00.00\pm 0$ 0.00 $0.20\pm 0.08$ $\mathbf{0.22\pm 0.06}$ $00.00\pm 0$ 0.00
@@
Door $k=0$ $0.08\pm 0.12$ $\mathbf{0.22\pm 0.06}$ $00.00\pm 0$ 0.00 $0.08\pm 0.12$ $\mathbf{0.22\pm 0.06}$ $00.00\pm 0$ 0.00
@@
Seen Categories $k=0$ $\mathbf{0.66\pm 0.09}$ $1.21\pm 0.06$ $1.45\pm 0.08$ $\mathbf{0.66\pm 0.09}$ $1.21\pm 0.06$ $1.45\pm 0.08$ $1.45\pm 0.08$
@@
Door $k=0$ $0.21\pm 0.04$ $\mathbf{0.22\pm 0.06}$ $00.00\pm 0$ 0.00 $0.21\pm 0.04$ $\mathbf{0.22\pm 0.06}$ $00.00\pm 0$ 0.00
@@
Door $k=0$ $0.20\pm 0.08$ $\mathbf{0.22\pm 0.06}$ $00.00\pm 0$ 0.00 $0.20\pm 0.08$ $\mathbf{0.22\pm 0.06}$ $00.00\pm 0$ 0.00
@@
Door $k=0$ $0.08\pm 0.12$ $\mathbf{0.22\pm 0.06}$ $00.00\pm 0$ 0.00 $0.08\pm 0.12$ $\mathbf{0.22\pm 0.06}$ $00.00\pm 0$ 0.00
@@
Seen Categories $k=0$ $\mathbf{0.66\pm 0.09}$ $1.21\pm 0.06$ $1.45\pm 0.08$ $\mathbf{0.66\pm 0.09}$ $1.21\pm 0.06$ $1.45\pm 0.08$